A body obeying simple harmonic motion has time period 2 seconds and amplitude 2 meters. Find maximum velocity and KE between mean position and extreme position if the mass of the body is 50 gm.

Grade:11

1 Answers

Mahima Kanawat
6 years ago
Dear student ,Vmax = aw = a [ 2Π/T ] = 2 ×[ 2Π/2 ]= 2Π KE = 1/2mv² = 1/2 × 1/20 × [4Π²] = Π²/10 Which is nearly equal to 1J With Regards MahimaAskiitians forum